OpenStudy (anonymous):

I'm having a really hard time with 2 questions, could you help and explain please? 1.Using complete sentences, explain how to graph y = 1/2x - 3 using the slope-intercept method. 2. Explain how to graph x + 3y = 9 using the x- and y-intercepts method. Be sure to use complete sentences and use the x- and y-intercepts method only

OpenStudy (anonymous):

for \(y=\frac{1}{2}x-3\) start with a point on \((0,-3)\) then since the slope is \(\frac{1}{2}\) go to the right two units and up one unit for example \((2,-2)\) is on the graph, also \((4,0)\) and \((6,1)\) etc

OpenStudy (anonymous):

for \(x+3y=9\) if \(x=0\) then you get \(3y=9\implies y=3\) so you know \((0,3)\) is on the graph also if \(y=0\) you get \(x=9\) so \((9,0)\) is on the graph plot the points, connect the dots.

OpenStudy (anonymous):

Ohhhhhhhhhhhhhhhhhh, you make it seem so easy. Thank you soooo much, this is so helpful